OpenStudy (anonymous):

Find the discount. regular price = $420; discount rate = 1/4 A. $1.05 B. $10.50 C. $105 D. $315

OpenStudy (nincompoop):

? this is incomplete

OpenStudy (beccaboo333):

I was gonna say that too.. Is there more to this problem?

OpenStudy (nincompoop):

think of a dollar and each quarter is 25 cents representing .25, which is 1/4 4 quarters means 4 * .25 = 1.00 now think of 420 and divide it into 4 420 divided by 2 is 210 and 210 divided by 2 is ______ that is a long method of analyzing the problem the easy way is to do this \[\frac{ 420 }{ 4 }\] think of how many 4's can you put in 400 and then how many 4's can you put in 20 add those together and that should be your answer

OpenStudy (anonymous):

regular price * discount rate= 420*1/4=105

OpenStudy (anonymous):

oh ok i get it thakn you

OpenStudy (nincompoop):

omg....

OpenStudy (anonymous):

o.0 wat?